# Adaptive Sliding Mode Control for Vehicle Platoons with State-Dependent Friction Uncertainty

### v1

@@ Abstract
Parts of this work were previously presented in [yadav2021adaptive]; the present manuscript provides a consolidated and expanded exposition with complete derivations and simulations
@@ 1.3.1 Motivation
Cooperative mobile multi-robot systems have gained significant attention due to their advantages over single-robot systems, such as better efficiency, high tolerance, redundancy, and manoeuvrability [darmanin2017review, arai2002advances]. These multi-robot systems have a broad range of applications, including search and rescue, exploration, navigation, security and surveillance, precision agriculture, and payload transportation [ismail2018survey, cai2012survey, cortes2017coordinated, soni2018formation]. One specific application is the platoon of autonomous vehicles, where multiple agents/robots/vehicles follow a common path in a shared environment. The control objective of such formation control is often decentralized [antonelli2006kinematic] , where the aim is to maintain a desired line-of-sight range between each vehicle and its predecessor while proceeding along a given trajectory [dai2017platoon]. However, controlling these systems under parametric uncertainties and unmodelled dynamics is still a significant challenge and an open problem.
@@ 1.3.2 Related Works
When it comes to autonomous platooning, the accuracy of formation control is heavily influenced by vehicle dynamics and external disturbances, particularly at higher speeds. The use of a simple Coulomb friction model is inadequate to capture the complex relationship between tyre wear and friction forces, as evidenced by previous studies [berntorp2019steering, iurian2005identification, vcerkala2014mobile]. However, accurately parameterizing this phenomenon in real-world situations is challenging, if not impossible, due to the variability of friction forces caused by factors such as changes in payload quantity, road conditions, and tyre distortion.
Advanced control strategies such as interpolating control [tuchner2017vehicle], distributed formation control [xu2019distributed], intervehicle distance control [zhang2020analysis], and robust control [zhao2018robust] have been developed to handle the dynamics of vehicles for both longitudinal and lateral control in automatic platoon formation. However, these strategies require prior knowledge of system parameters. To address this limitation, two-stage tracking controllers have been used for wheeled mobile robots [chen2009design, asif2014adaptive, koubaa2018adaptive], combining a kinematic controller with an adaptive sliding mode controller (ASMC). However, these controllers can only handle bounded uncertainties, and frictional forces, which are state-dependent, do not satisfy this requirement [roy2019overcoming, utkin2013adaptive].
@@ 2.2 CONTROL FORMULATION
Controlling a nonholonomic WMR effectively requires considering both kinematic and dynamic model-based controllers [chen2009design, asif2014adaptive, koubaa2018adaptive]. In this approach, the linear and angular velocity derived from the kinematic controller are used as the desired trajectory in the dynamic model. However, since parametric uncertainty can only be captured in the dynamic model, the novelty of this work lies in designing an ASMC for the dynamic model while using the standard kinematic controller as in previous works [fierro1997control, garcia2017tracking, liu2020adaptive].
@@ 2.2.1 Kinematic Control Design
Following [fierro1997control, garcia2017tracking, liu2020adaptive], we used the following backstepping method based kinematic tracking control law
@@ 2.2.2 PROPOSED DYNAMIC CONTROLLER DESIGN
It is crucial to note that viscous friction being proportional to the velocity $v$ of the vehicle, the friction forces are state-dependent and thus, cannot be bounded a priori [roy2020towards, roy2020adaptive, roy2021adaptive, shukla2021robust]. Such consideration segregates this work from the state-of-the-art adaptive solutions [chen2009design, asif2014adaptive, koubaa2018adaptive] relying on a priori bounded dynamical forces.
@@ 2.3 Stability Analysis of The Proposed Controller
To avoid chattering due to discontinuity in control law, the ‘sgnsgn’ functions in (2.11) and (2.18) are often replaced by a ‘saturation’/sigmoid functions which leads to minor modifications in the stability analysis without altering the overall UUB result and hence omitted to avoid repetition (cf. [roy2019overcoming, asif2014adaptive, roy2020new]).
@@ 3.0.1 Simulation Scenario
The study aims to compare the performance of the proposed adaptive sliding mode controller (ASMC) with the standard ASMC [asif2014adaptive]. The evaluation is conducted in a Gazebo simulation platform using an open-source TurtleBot3 robot model [turtlebot] and a Teeterbot plugin [teeterbot] to give a torque command to each motor using (1.11). The controller’s performance is assessed in a custom arena in Gazebo  [^1] Simulation Video: https://youtu.be/7yHY9atSeK8; Titled as ”Adaptive Sliding Mode Control for Autonomous Vehicle Platoon under Unknown Friction Forces” on Youtube, which is divided into four quadrants with different surface longitudinal friction values. The first, second, and fourth quadrants have a friction value of $\mu_{1}=0.1$, while the third quadrant has a friction value of $\mu_{2}=0.13$. The lateral friction value is kept constant at $0.1$ in all quadrants to prevent lateral slippage of the robots. Additionally, two speed breakers are placed in the arena to provide sudden interruptions (cf. Fig. 3.1).
@@ 3.0.2 Parameter Selection
For simulation, the kinematic control parameters are selected as: $k_{1}$ = 5, $k_{2}$ = 3, $k_{3}$ = 2, $v_{d}=2$ m/sec for all robots. The control parameters of the proposed ASMC are selected to be: $\phi_{v}=0.5,\phi_{w}=0.1,\Lambda_{v}=3,\Lambda_{w}=2,\hat{K}_{v0}(0)=\hat{K}_{v1}(0)=\hat{K}_{w2}(0)=\hat{K}_{w0}(0)=\hat{K}_{w1}(0)=\hat{K}_{v2}(0)=0.01,\alpha_{v0}=2.5,\alpha_{v1}=2.5,\alpha_{w2}=3,\alpha_{w0}=5,\alpha_{w1}=5,\alpha_{v2}=1.5$ for all robots. For parity, similar sliding variable and similar kinematic control parameters are designed for the standard ASMC [asif2014adaptive].
@@ 3.0.3 Results and Analysis
The study compared the performance of two controllers, the proposed Adaptive Sliding Mode Controller (ASMC) and the standard ASMC [asif2014adaptive] via Figs. 3.3, 3.4 and 3.5, in terms of position tracking error for three robots, labeled 1, 2, and 3. The analysis was carried out in four quadrants. In the third quadrant, between $300<t<400$, all robots experienced a sharp turn and high friction surface, causing an increase in tracking error. The results show that the standard ASMC [asif2014adaptive] performed worse than the proposed ASMC during this period, as it was not designed to handle changes in state-dependent friction components. A similar trend was observed between $500<t<600$, when the robots encountered two speed breakers. The proposed controller showed fewer spikes in the tracking error profile during this period. The performance of both controllers was found to be similar in the first, second, and fourth quadrants. This was confirmed by path tracking performances shown in Figures 3.6, 3.7 and 3.8. Overall, the results indicate that the proposed controller performs better in the third quadrant, where changes in friction components are significant, while the standard and proposed controllers have similar performance in the other quadrants.
@@
Additionally, in Figure 3.9, the difference between the intended gap and the actual gap of two consecutive robots, namely Robot 1-Robot 2 and Robot 2-Robot 3, is illustrated, which is called the gap error. At around $t=350$ seconds, it can be observed that when the robots are in the third quadrant and encounter a high variation in friction, the standard ASMC [asif2014adaptive] exhibits a significantly greater gap error than the proposed ASMC. To provide more conclusive evidence, Tables 3.1 and 3.2 present the performance of both controllers in terms of root mean-squared (RMS) error. The data in these tables indicate that the proposed ASMC achieves superior tracking accuracy while preserving the desired distance between the robots.
@@
    Adaptive Sliding Mode Control for Autonomous Vehicle Platoon under Unknown Friction Forces [yadav2021adaptive]
@@
    Robustifying Payload Carrying Operations for Quadrotors Under Time-Varying State Constraints and Uncertainty [sankaranarayanan2022robustifying]
@@
  2. 2.
    Efficient Manoeuvring of Quadrotor under Constrained Space and Predefined Accuracy [ganguly2021efficient]
  3. 3.
    CCO-VOXEL: Chance Constrained Optimization over Uncertain Voxel-Grid Representation for Safe Trajectory Planning [harithas2022cco]
@@
  4. 4.
    Robust Manoeuvring of Quadrotor under Full State Constraints [ganguly2022robust]
    Automatic Control and Dynamical Optimization Society (ACDOS 2022)
@@
    Introducing Scissor Mechanism based Novel Reconfigurable Quadrotor: Design, Modelling and Control [suraj2022introducing]
@@
    Adaptive Artificial Time Delay Control for Quadrotors under State-dependent Unknown Dynamic [dantu2022adaptive]
@@
    Adaptive Anti-swing Control for Clasping Operations in Quadrotors with Cable-suspended Payload [dantu2023adaptive]
@@
    An Integrated Approach to Aerial Grasping: Combining a Bistable Gripper with Adaptive Control [yadav2023integratedapproachaerialgrasping]

### v2

@@ Abstract
Parts of this work were previously presented in [41]; the present manuscript provides a consolidated and expanded exposition with complete derivations and simulations
@@ 1.3.1 Motivation
Cooperative mobile multi-robot systems have gained significant attention due to their advantages over single-robot systems, such as better efficiency, high tolerance, redundancy, and manoeuvrability [13, 2]. These multi-robot systems have a broad range of applications, including search and rescue, exploration, navigation, security and surveillance, precision agriculture, and payload transportation [20, 5, 8, 34]. One specific application is the platoon of autonomous vehicles, where multiple agents/robots/vehicles follow a common path in a shared environment. The control objective of such formation control is often decentralized [1] , where the aim is to maintain a desired line-of-sight range between each vehicle and its predecessor while proceeding along a given trajectory [9]. However, controlling these systems under parametric uncertainties and unmodelled dynamics is still a significant challenge and an open problem.
@@ 1.3.2 Related Works
When it comes to autonomous platooning, the accuracy of formation control is heavily influenced by vehicle dynamics and external disturbances, particularly at higher speeds. The use of a simple Coulomb friction model is inadequate to capture the complex relationship between tyre wear and friction forces, as evidenced by previous studies [4, 21, 6]. However, accurately parameterizing this phenomenon in real-world situations is challenging, if not impossible, due to the variability of friction forces caused by factors such as changes in payload quantity, road conditions, and tyre distortion.
Modern robotics research increasingly emphasizes reliable operation under uncertainties, constraints, and complex system dynamics. To address these challenges, researchers have explored a range of methods including adaptive control, robust system design, and advanced planning strategies. Such approaches play a crucial role in improving the performance and dependability of robotic systems in practical applications. Representative studies in this area are presented in [15, 16, 31, 19, 35, 12, 10, 18, 39, 11, 40, 32].
Advanced control strategies such as interpolating control [36], distributed formation control [38], intervehicle distance control [42], and robust control [43] have been developed to handle the dynamics of vehicles for both longitudinal and lateral control in automatic platoon formation. However, these strategies require prior knowledge of system parameters. To address this limitation, two-stage tracking controllers have been used for wheeled mobile robots [7, 3, 22], combining a kinematic controller with an adaptive sliding mode controller (ASMC). However, these controllers can only handle bounded uncertainties, and frictional forces, which are state-dependent, do not satisfy this requirement [30, 37].
@@ 2.2 CONTROL FORMULATION
Controlling a nonholonomic WMR effectively requires considering both kinematic and dynamic model-based controllers [7, 3, 22]. In this approach, the linear and angular velocity derived from the kinematic controller are used as the desired trajectory in the dynamic model. However, since parametric uncertainty can only be captured in the dynamic model, the novelty of this work lies in designing an ASMC for the dynamic model while using the standard kinematic controller as in previous works [14, 17, 23].
@@ 2.2.1 Kinematic Control Design
Following [14, 17, 23], we used the following backstepping method based kinematic tracking control law
@@ 2.2.2 PROPOSED DYNAMIC CONTROLLER DESIGN
It is crucial to note that viscous friction being proportional to the velocity $v$ of the vehicle, the friction forces are state-dependent and thus, cannot be bounded a priori [28, 26, 27, 33]. Such consideration segregates this work from the state-of-the-art adaptive solutions [7, 3, 22] relying on a priori bounded dynamical forces.
@@ 2.3 Stability Analysis of The Proposed Controller
To avoid chattering due to discontinuity in control law, the ‘sgnsgn’ functions in (2.11) and (2.18) are often replaced by a ‘saturation’/sigmoid functions which leads to minor modifications in the stability analysis without altering the overall UUB result and hence omitted to avoid repetition (cf. [30, 3, 29]).
@@ 3.0.1 Simulation Scenario
The study aims to compare the performance of the proposed adaptive sliding mode controller (ASMC) with the standard ASMC [3]. The evaluation is conducted in a Gazebo simulation platform using an open-source TurtleBot3 robot model [24] and a Teeterbot plugin [25] to give a torque command to each motor using (1.11). The controller’s performance is assessed in a custom arena in Gazebo  [^1] Simulation Video: https://youtu.be/7yHY9atSeK8; Titled as ”Adaptive Sliding Mode Control for Autonomous Vehicle Platoon under Unknown Friction Forces” on Youtube, which is divided into four quadrants with different surface longitudinal friction values. The first, second, and fourth quadrants have a friction value of $\mu_{1}=0.1$, while the third quadrant has a friction value of $\mu_{2}=0.13$. The lateral friction value is kept constant at $0.1$ in all quadrants to prevent lateral slippage of the robots. Additionally, two speed breakers are placed in the arena to provide sudden interruptions (cf. Fig. 3.1).
@@ 3.0.2 Parameter Selection
For simulation, the kinematic control parameters are selected as: $k_{1}$ = 5, $k_{2}$ = 3, $k_{3}$ = 2, $v_{d}=2$ m/sec for all robots. The control parameters of the proposed ASMC are selected to be: $\phi_{v}=0.5,\phi_{w}=0.1,\Lambda_{v}=3,\Lambda_{w}=2,\hat{K}_{v0}(0)=\hat{K}_{v1}(0)=\hat{K}_{w2}(0)=\hat{K}_{w0}(0)=\hat{K}_{w1}(0)=\hat{K}_{v2}(0)=0.01,\alpha_{v0}=2.5,\alpha_{v1}=2.5,\alpha_{w2}=3,\alpha_{w0}=5,\alpha_{w1}=5,\alpha_{v2}=1.5$ for all robots. For parity, similar sliding variable and similar kinematic control parameters are designed for the standard ASMC [3].
@@ 3.0.3 Results and Analysis
The study compared the performance of two controllers, the proposed Adaptive Sliding Mode Controller (ASMC) and the standard ASMC [3] via Figs. 3.3, 3.4 and 3.5, in terms of position tracking error for three robots, labeled 1, 2, and 3. The analysis was carried out in four quadrants. In the third quadrant, between $300<t<400$, all robots experienced a sharp turn and high friction surface, causing an increase in tracking error. The results show that the standard ASMC [3] performed worse than the proposed ASMC during this period, as it was not designed to handle changes in state-dependent friction components. A similar trend was observed between $500<t<600$, when the robots encountered two speed breakers. The proposed controller showed fewer spikes in the tracking error profile during this period. The performance of both controllers was found to be similar in the first, second, and fourth quadrants. This was confirmed by path tracking performances shown in Figures 3.6, 3.7 and 3.8. Overall, the results indicate that the proposed controller performs better in the third quadrant, where changes in friction components are significant, while the standard and proposed controllers have similar performance in the other quadrants.
@@
Additionally, in Figure 3.9, the difference between the intended gap and the actual gap of two consecutive robots, namely Robot 1-Robot 2 and Robot 2-Robot 3, is illustrated, which is called the gap error. At around $t=350$ seconds, it can be observed that when the robots are in the third quadrant and encounter a high variation in friction, the standard ASMC [3] exhibits a significantly greater gap error than the proposed ASMC. To provide more conclusive evidence, Tables 3.1 and 3.2 present the performance of both controllers in terms of root mean-squared (RMS) error. The data in these tables indicate that the proposed ASMC achieves superior tracking accuracy while preserving the desired distance between the robots.
@@
    Adaptive Sliding Mode Control for Autonomous Vehicle Platoon under Unknown Friction Forces [41]
@@
    Efficient Manoeuvring of Quadrotor under Constrained Space and Predefined Accuracy [15]
  2. 2.
    Robust Manoeuvring of Quadrotor under Full State Constraints [16]
    Automatic Control and Dynamical Optimization Society (ACDOS 2022)
  3. 3.
    Robustifying Payload Carrying Operations for Quadrotors Under Time-Varying State Constraints and Uncertainty [31]
@@
  4. 4.
    CCO-VOXEL: Chance Constrained Optimization over Uncertain Voxel-Grid Representation for Safe Trajectory Planning [19]
@@
    Introducing Scissor Mechanism based Novel Reconfigurable Quadrotor: Design, Modelling and Control [35]
@@
    Adaptive Artificial Time Delay Control for Quadrotors under State-dependent Unknown Dynamic [12]
@@
    Adaptive Anti-swing Control for Clasping Operations in Quadrotors with Cable-suspended Payload [10]
@@
    Adaptive Control of Quadrotor under Actuator Loss and Unknown State-dependent Dynamics [18]
    Saksham Gupta, Amitabh Sharma, Aditya Mulgundkar, Rishabh Dev Yadav, Spandan Roy
    IEEE International Conference on Automation Science and Engineering (CASE)
  9. 9.
    Modular adaptive aerial manipulation under unknown dynamic coupling forces [39]
    Rishabh Dev Yadav, Swati Dantu, Wei Pan, Sihao Sun, Spandan Roy, Simone Baldi
    IEEE/ASME Transactions on Mechatronics
  10. 10.
    Adaptive Tracking and Anti-Swing Control of Quadrotors Carrying Suspended Payload Under State-Dependent Uncertainty [11]
    Swati Dantu, Rishabh Dev Yadav, Ananth Rachakonda, Spandan Roy, Simone Baldi
    IEEE/ASME Transactions on Mechatronics
  11. 11.
    An Integrated Approach to Aerial Grasping: Combining a Bistable Gripper with Adaptive Control [40]
@@
  12. 12.
    Impedance and Stability Targeted Adaptation for Aerial Manipulator with Unknown Coupling Dynamics [32]
    Amitabh Sharma, Saksham Gupta, Shivansh Pratap Singh, Rishabh Dev Yadav, Hongyu Song, Wei Pan, Spandan Roy, Simone Baldi
    IEEE International Conference on Control, Automation and Systems (ICCAS)